Hello, Could you help me with this equation: Solve and correct the error. x^3-15x5=4+2x^2+1 x^3-2x^2-15x=0 x(x^2-2x-15)=0 (x-5)(x+3)=0 x={-3,5}
Answer by ikleyn(52787)

.
Hello, Could you help me with this equation: Solve and correct the error.
x^3-15x5=4+2x^2+1 x^3-2x^2-15x=0 x(x^2-2x-15)=0 (x-5)(x+3)=0 x={-3,5}
--------------------------------------------------------------------------

How to read this term, -15x5, the second term in the left side of the first equation ??

What is it ??